A juggler tosses three balls alternately a vertically upward. Each ball has an initial velocity of 5 m/s. How high does each ball rise? How long does each ball remain in the air?

1
Expert's answer
2021-12-20T10:31:47-0500
"h=\frac{v^2}{2g}\\h=\frac{5^2}{2(9.8)}=1.3\ m\\t=2\frac{5}{9.8}=1.0\ s"